The length of a rectangle is 5 times its width. If its area is 125 square feet, what is its dimensions?
Answer by jack.wang5(5)

Let A = area = L W, W = width, L = length = 5W.
Therefore A = L W = 5 W W = 5 W^2 , so A = 5 W^2.
Solve W , and thus W = sqrt(A/5).
Since A = 125, therefore W = sqrt(125/5) = 5.